# CliniBench: A Clinical Outcome Prediction Benchmark for Generative and Encoder-Based Language Models

### v1

@@ Abstract
With their growing capabilities, generative large language models (LLMs) are being increasingly investigated for complex medical tasks. However, their effectiveness in real-world clinical applications remains underexplored. To address this, we present CliniBench, the first benchmark that enables comparability of well-studied encoder-based classifiers and generative LLMs for discharge diagnosis prediction from admission notes in MIMIC-IV dataset. Our extensive study compares 12 generative LLMs and 3 encoder-based classifiers and demonstrates that encoder-based classifiers consistently outperform generative models in diagnosis prediction. We assess several retrieval augmentation strategies for in-context learning from similar patients and find that they provide notable performance improvements for generative LLMs.
@@ 1 Introduction
Encoder-based classifiers
@@
  • We publish the first benchmark [^1] URL to the benchmark anonymized during the review. We release the benchmark, admission IDs, and a script for dataset construction. The clinical notes are not shared due to the restrictions of the PhysioNet license. By publishing our scripts to build the dataset and splits based on the admission IDs, we guarantee full reproducibility of our results. to compare encoders with generative LLMs for diagnosis prediction from admission notes based on MIMIC-IV Johnson et al. (2023c).
@@ 2 Related Work
Clinical outcome prediction is based solely on information available at hospital admission represents a distinct challenge compared to retrospective ICD coding or medical question answering tasks. van Aken et al. (2021) introduce the task based on MIMIC-III Goldberger et al. (2000); Johnson et al. (2016) to forecast discharge outcomes such as ICD codes, mortality, and length-of-stay—based. Naik et al. (2022) propose enhancing outcome prediction by retrieving relevant biomedical literature, grounding the predictions to improve both accuracy and interpretability. Röhr et al. (2024) extend this task to MIMIC-IV Johnson et al. (2023c), adding new outcomes like patient routing and evaluating benchmark performance. Grundmann et al. (2024) examine data drift between MIMIC versions and highlighted inconsistencies in ICD documentation practices, taxonomy, and label reliability. To improve label generalization and interpretability, van Aken et al. (2022) propose ProtoPatient, a prototypical network-based model. This approach was later extended by Figueroa et al. (2024) to better handle long-tail ICD distributions. Gema et al. (2024) demonstrate that fine-tuning LLMs for classification, rather than generation, can yield improved performance on admission-based tasks using parameter-efficient methods. Shoham and Rappoport (2024) focus on predicting readmission for three specific conditions, using LLMs trained on structured diagnosis descriptions from prior visits. In contrast to similar studies that utilize the full patient notes as inputs, the authors use textual descriptions of previously assigned diagnostic codes for their prediction.
@@ 2.1 Related Tasks
Two common well studied, related tasks are ICD coding and medical question answering. Furthermore, as noted by Bedi et al. (2025), only five percent currently make use of real patient records and realistic clinical tasks.
@@ 2.2 Learning at Inference Time
enhances LLMs by incorporating external knowledge via in-context-learning, improving accuracy, and reducing hallucinations. For example, Lu et al. (2024) introduced ClinicalRAG for reliable clinical decision support, Kresevic et al. (2024) used RAG to increase guideline interpretation accuracy to 99% and Unlu et al. (2024) applied RAG for efficient clinical trial screening, outperforming manual methods. While Xiong et al. (2024) benchmarked RAG systems, showing up to 18% accuracy improvements, Ngo et al. (2024) provided practical optimization guidelines for RAG systems. Finally, Li et al. (2025) show that chunking and fine-granular data selection for RAG, improves performance on biomedical relation extraction and classification tasks.
@@ 3 CliniBench
In order to compare encoder-based classifiers and generative LLMs we build the CliniBench benchmark depicted in Fig.˜1. The benchmark consists of the task to predict a patients outcome in the form of discharge codes for a medical text.
@@ 3.1 MIMIC-IV Dataset
The source of medical texts is the dataset Medical Information Mart for Intensive Care (MIMIC-IV). MIMIC-IV is grouped into two modules: hosp and icu. Data in the hosp module are sourced from the hospital-wide electronic health record (EHR), while data in the icu module are sourced from the clinical information system from the intensive care unit (ICU).
@@
To support clinical decision-making through diagnosis prediction, we limit the extracted information to relevant data known at admission time: Chief Complaint, Major Surgical or Invasive Procedure, Allergies, History of Present Illness, Past Medical History, Social History, Family History, Physical Exam at Admission, and Medication at Admission.
To avoid data leakage and maintain temporal consistency, we exclude sections containing outcome-related information obtained during the stay or after the course of hospitalization: Physical Exam during Stay and at Discharge, Pertinent Results, Brief Hospital Course, Medication at Discharge, Discharge Disposition, Facility, Discharge Diagnosis, Discharge Condition, Discharge Instructions, and Followup Instructions.
Finally, we exclude any administrative or demographic fields that are either de-identified or do not contribute to the diagnosis prediction task: Name, Unit No, Admission Date, Discharge Date, Date of Birth, Sex, Service, and Attending.
@@
Both modules include detailed discharge codes, derived from the International Classification of Diseases (ICD) coding system. This documents diseases and medical conditions (morbidity) associated with each admission. Depending on the date of admission, the codes are available either in ICD-9 or in the newer ICD-10 formats (see Table˜1). Given the hierarchical nature of the ICD system, we reduce the granularity of each code to three digits, allowing encoder-based models to learn from multiple examples rather than just a single instance during fine-tuning. This reduces the average number of codes per note to 13.98 while still providing adequate clinical information.
@@
For a more comprehensive error analysis, we divide the data by ICD version and module. This is due to the differences of the distribution of codes between icu and hosp settings. Detailed information about the splits can be found in Table˜1.
@@
#Notes 38708 26007 170599 93158
@@
Avg. #Codes/Note 14.6 18.4 10.0 12.9
Avg. #Tokens/Note 670.4 624.9 681.5 687.5
Threshold, Tail 8 7 17 10
Threshold, Head 55 43 126 67
Table 1: Statistics of the four datasets derived from MIMIC-IV. The number of tokens is derived by using the BioMedBERT tokenizer. Below we report the absolute tertile thresholds for head (1st tertile) and tail (3rd tertile). This means that each of the diagnosis codes in the tail tertile appears less than 8 times (icu ICD-9 dataset)
@@ 3.2 Model Classes and Augmentation
For the encoder-based prediction pipeline, we fine-tune a pretrained transformer to predict ICD diagnosis code probabilities. To ensure comparability with generative models, we evaluate using threshold-based metrics like recall and precision. Since thresholds can be noisy for rare classes, we tune them by maximizing the $F_{1}$ score on the Precision-Recall curve. We report details on the threshold-tuning in the Appendix in Fig.˜6.
@@
For the generative LLMs, we prompt each model to output a list of diagnoses short descriptions. We apply guided decoding Willard and Louf (2023) to ensure a machine readable JSON format. These structured outputs are mapped to corresponding ICD codes using text-based matching. We investigate three distinct approaches: Zero-Shot, Few-Shot, and Zero-Shot with CoT prompting.
@@
In the Zero-Shot setting, we input the admission notes with task specific instructions. Thus, the model can only leverage knowledge acquired from its initial pretraining and the information available in the shown example.
@@
Generative LLMs can learn from few-shot examples provided in the prompt Brown et al. (2020); Song et al. (2023); Parnami and Lee (2022) without the adaption of parameters (in-context-learning). For this adaption, either the user or a system incorporates complementary knowledge in the prompt. This is especially useful for domain specific tasks that are not well represented in the pretraining data.
@@ 4 Experimental Setup
We evaluate the selected models on the top-$n$ predicted diagnoses using macro averaged recall, precision, and $F_{1}$ score. The evaluation of the top-$n$ diagnoses is necessary to ensure comparability between encoder-based models and the fixed number of predictions produced by the generative LLMs. As a trade-off between the average number of codes in each dataset split and to provide headroom for erroneous predictions, we choose $n=20$. Furthermore, limiting the number of codes increases computational efficiency due to a less complex regular expression for the guided generation and shorter generated sequences. Additionally, we report mean average precision (MAP), which does not rely on thresholding. Finally, we assess main diagnosis accuracy (MD Acc.), which indicates whether the first annotated label (main diagnosis) is included in the predicted diagnoses.
@@ 4.1 Encoder
We fine-tune the pretrained encoders on the train split and select the best performing using AUROC on the validation split using AdamW Loshchilov and Hutter (2019) as optimizer and a polynomial learning rate decay. For the hyperparameter optimization, we apply Bayesian optimization. For more details compare Table˜6.
@@ 4.2 Generative LLMs
We evaluate open weight generative LLMs that are capable of processing sequences exceeding 4096 tokens. This extended sequence capacity is crucial for evaluating models in the few-shot setting, as sequence length increases with each added demonstration. Consequently, we excluded both available medical LLMs MedAlpaca and Med42 from our analysis, as both are based on the Llama2 architecture, which limits their maximum sequence length to 2048 tokens. We selected the following models for evaluation: Llama 3.1 8B and 70B, Mistral 7B, Mistral Nemo (13B), Qwen2 7B and 72B. For each model, we evaluate both the instruction fine-tuned and the non-instruction-tuned variants, resulting in a total of 12 models. We explicitly do not benchmark closed-source commercial models due to constraints of the Physionet license  [^2] https://physionet.org/content/mimiciv/view-license/2.2/.
@@
While generative models often produce valid JSON outputs and the correct number of prompted diagnoses, some of the evaluated models, especially those not fine-tuned with instruction data, frequently fail to do so. This limitation reduces overall task performance compared to encoder-based models. We observe that smaller models tend to produce valid JSON less reliably (e.g., Qwen2 7B Instruct 97.26% valid JSON vs. Qwen2 72B 100%; Llama 3.1 8B Instruct 0.49% vs. Llama 3.1 70B Instruct 3.82%). None of the non-instruction-tuned models are capable of generating valid JSON. We report the performance of generating valid JSON output for all models in the Appendix in Table˜4. To mitigate this issue, we apply guided decoding Willard and Louf (2023) to enforce JSON output that contains exactly 20 strings. Each string represents a diagnosis description between 3 and 70 characters in length. To optimize performance, we limit the number of generated tokens to a maximum of 1500. Finally, we use greedy sampling with a temperature of 0.
@@
We use multiple prompts for each experiment: Zero-shot, zero-shot with CoT, and few-shot learning. The prompts consist each of a role description for the model and the respective task. For few-shot demonstrations, we provide the admission notes together with their annotated diagnosis short descriptions in the desired JSON format. We report the used prompts in Section˜A.1 We hypothesize that providing the model with comprehensive information about the ICD ontology would enhance its performance. However, the full set of code descriptions exceeds the maximum context length of all evaluated models. Using the Qwen2.5 tokenizer, the descriptions of all possible ICD-9 codes require 167,679 tokens, and those of ICD-10 codes require 1,369,475 tokens. Consequently, we restrict the input to the task prompt, admission note and few-shot examples.
@@
Because LLMs often struggle to predict exact ICD codes Boyle et al. (2023), we instead generate diagnostic descriptions and map them to ICD codes via their corresponding long or short descriptions. While semantic matching is guaranteed to resolve all ICD descriptions, BM25 discards 14.2% of the predicted diagnosis descriptions over all experiments. Based on a comparison of different sentence transformers, we adopt NeuML-PubMedBERT [^3] https://huggingface.co/NeuML/pubmedbert-base-embeddings for text-to-label mapping. Details are reported in the Appendix in Table˜5. Finally, we truncate ICD codes to three digits to reduce label space and remove duplicates from the predictions.
CoT implementation.
@@ 4.3 Retrieval Models and Few-Shot Learning
We implement few-shot learning by incorporating one, three, or five demonstrations into the prompt.
The effectiveness of in-context-learning depends on the selection of demonstrations Liu et al. (2022). Thus, we use a retriever to find the most relevant candidates. The retriever takes one admission note as input query and returns the most similar admission notes from the training split.
We conduct three different retrieval strategies and compare them to random and gold demonstrations. Namely, we use document-based semantic similarity, latent outcome similarity (classifier fine-tuned on diagnosis prediction task), and a term-based retrieving strategy.
@@
To assess the performance of retrieval models for clinical outcome prediction, we evaluate them directly on the diagnosis prediction task. This enables us to gauge the impact of adding a generative LLM on top of the retrieved few-shot results. We apply majority voting over the annotated labels of five retrieved documents from the training split. From these documents, we subsequently select the top 20 most frequent labels. Labels from documents with a higher document similarity are prioritized in case of equal frequency. We evaluate the following semantic similarity retrievers based on the S-BERT architecture Reimers and Gurevych (2019): NeuML-PubmedBERT, jina-base-v2 Günther et al. (2023), all-mpnet-base-v2444https://huggingface.co/sentence-transformers/all-mpnet-base-v2, all-mpnet-base-v2-neg Anschütz et al. (2023), BiomedBERT Gu et al. (2021), and S-PubMedBert-MS-MARCO Deka et al. (2022). To reduce computational costs in the few-shot experiments, we choose only the best-performing model from the semantic similarity retrievers. For the latent outcome similarity retrieval, we employ our fine-tuned baseline encoder (BioMedBERT) as specified in Section˜4. We use the CLS token from the last embedding layer as a representation of the admission note. We expect that admission notes with similar outcomes cluster together in the embedding space and consequently outperform semantic similarity. Finally, we employ BM25 Robertson and Zaragoza (2009) for term-based retrieval.

### v2

@@ Abstract
With their growing capabilities, generative large language models (LLMs) are being increasingly investigated for complex medical tasks. However, their effectiveness in real-world clinical applications remains underexplored. To address this, we present CliniBench, the first benchmark that enables comparability of well-studied encoder-based classifiers and generative LLMs for discharge diagnosis prediction from admission notes in the MIMIC-IV dataset. Our extensive study compares 12 generative LLMs and 3 encoder-based classifiers and demonstrates that encoder-based classifiers consistently outperform generative models in diagnosis prediction. We assess several retrieval augmentation strategies for in-context learning from similar patients and find that they provide notable performance improvements for generative LLMs.
CliniBench: A Clinical Outcome Prediction Benchmark for Generative and Encoder-Based Language Models
Paul Grundmann1, Jan Frick1, Dennis Fast1, Thomas Steffek1, Felix Gers1, Wolfgang Nejdl2, Alexander Löser1, 1DATEXIS, Berlin University of Applied Sciences, Berlin, Germany 2Leibniz University Hannover, Hannover, Germany Correspondence: jan.frick@bht-berlin.de
@@ 1 Introduction
Encoder-based classifiers (Encoder)
@@
  • We publish the first benchmark [^1] https://github.com/datexis/clinibench to compare encoders with generative LLMs for diagnosis prediction from admission notes based on MIMIC-IV Johnson et al. (2023b).
@@ 2 Related Work
Clinical outcome prediction is based solely on information available at hospital admission represents a distinct challenge compared to retrospective ICD coding or predefined text classification tasks. van Aken et al. (2021) introduce the task based on MIMIC-III Goldberger et al. (2000); Johnson et al. (2016) to forecast discharge outcomes such as ICD codes, mortality, and length-of-stay. Naik et al. (2022) propose enhancing outcome prediction by retrieving relevant biomedical literature, grounding the predictions to improve both accuracy and interpretability. Röhr et al. (2024) extend this task to MIMIC-IV Johnson et al. (2023b), adding new outcomes like patient routing and evaluating benchmark performance. Grundmann et al. (2024) examine data drift between MIMIC versions and highlight inconsistencies in ICD documentation practices, taxonomy, and label reliability. To improve label generalization and interpretability, van Aken et al. (2022) propose ProtoPatient, a prototypical network-based model. This approach was later extended by Figueroa et al. (2024) to better handle long-tail ICD distributions. Gema et al. (2024) demonstrate that fine-tuning LLMs for classification, rather than generation, can yield improved performance on admission-based tasks using parameter-efficient methods. Shoham and Rappoport (2024) focus on predicting readmission for three specific conditions, using LLMs trained on structured diagnosis descriptions from prior visits. In contrast to similar studies that utilize the full patient notes as inputs, the authors use textual descriptions of previously assigned diagnostic codes for their prediction.
@@ 2.1 Related Tasks
Two common well-studied, related tasks are ICD coding and medical question answering. Furthermore, as noted by Bedi et al. (2025), only five percent currently make use of real patient records and realistic clinical tasks.
@@ 2.2 Learning at Inference Time
enhances LLMs by incorporating external knowledge via in-context learning, improving accuracy, and reducing hallucinations. For example, Lu et al. (2024) introduced ClinicalRAG for reliable clinical decision support, Kresevic et al. (2024) used RAG to increase guideline interpretation accuracy to 99% and Unlu et al. (2024) applied RAG for efficient clinical trial screening, outperforming manual methods. While Xiong et al. (2024) benchmarked RAG systems, showing up to 18% accuracy improvements, Ngo et al. (2024) provided practical optimization guidelines for RAG systems. Finally, Li et al. (2025) show that chunking and fine-granular data selection for RAG improve performance on biomedical relation extraction and classification tasks.
@@ 3 CliniBench
In order to compare encoder-based classifiers and generative LLMs we build the CliniBench benchmark depicted in Fig.˜1. The benchmark consists of the task of predicting a patient’s outcome in the form of discharge codes for a clinical note at admission.
@@ 3.1 MIMIC-IV Dataset
The source of medical texts is the Medical Information Mart for Intensive Care (MIMIC-IV) dataset. MIMIC-IV is grouped into two modules: HOSP and ICU. Data in the HOSP module are sourced from the hospital-wide electronic health record (EHR), while data in the ICU module are sourced from the clinical information system from the intensive care unit (ICU).
@@
To support clinical decision-making through diagnosis prediction, we limit the extracted information to relevant data known at admission time. We exclude sections containing outcome-related information obtained during the stay or after hospitalization, and exclude administrative or demographic fields that are de-identified or do not contribute to the diagnosis prediction task (see Section˜A.2 for complete section details).
@@
Both modules include detailed discharge codes derived from the International Classification of Diseases (ICD) coding system, documenting diseases and medical conditions (morbidity) associated with each admission. Depending on the date of admission, codes are available in either ICD-9 or ICD-10 formats (see Table˜1). Given the hierarchical nature of the ICD system, we reduce the granularity of each code to three digits, allowing encoder-based models to learn from multiple examples rather than single instances during fine-tuning. This reduces the average number of codes per note to 13.98 while still providing adequate clinical information.
Dataset Construction.
We derive CliniBench dataset from the MIMIC-IV-Note dataset Johnson et al. (2023c) and ICD discharge codes from the MIMIC-IV dataset Johnson et al. (2023a) following van Aken et al. (2021). Each instance corresponds to one hospital admission note (see Appendix A.2 for construction details).
@@
We generate a stratified split into training (70%), validation (10%), and test (20%) sets. The dataset comprises four partitions reflecting coding system and care setting: ICD-9/ICD-10 × HOSP/ICU. We exclude labels not present across all three splits from evaluation. We additionally report the notes-per-label cutoffs used to define frequency tertiles: Tail contains labels with frequency $\leq t_{tail}$, Head contains labels with frequency $t_{head}$ and Body contains labels with $t_{tail}>freq\leq t_{head}$. Cutoffs are computed on the train split per partition. Detailed dataset statistics are provided in Table 1.
@@
#Train Adm. Notes 27007 18596 119025 66736
#Val Adm. Notes 3618 2399 15683 8912
#Test Adm. Notes 8089 5600 35930 20058
@@
Avg. #Tokens/Adm. Note 670.4 624.9 681.5 687.5
Avg. #Codes/Adm. Note 14.6 18.4 10.0 12.9
#Adm. Notes/Code (Tail) $\leq$ 8 7 17 10
#Adm. Notes/Code (Head) >> 55 43 126 67
Table 1: Statistics of the four CliniBench partitions derived from MIMIC-IV admission notes. We report train/validation/test sizes, label set sizes, average label and token counts, and admission notes-per-label cutoffs for the frequency tertiles. Token counts are computed using the BioMedBERT tokenizer.
@@ 3.2 Model Classes and Augmentation
For the encoder-based prediction pipeline, we fine-tune a pretrained transformer to predict ICD diagnosis code probabilities. To ensure comparability with generative models, we evaluate using threshold-based metrics like recall and precision. Since thresholds can be noisy for rare classes, we tune them by maximizing the $F_{1}$ score on the Precision-Recall curve. We report details on the threshold-tuning in Section 4.1.
@@
For the generative LLMs, we prompt each model to output a list of short diagnosis descriptions. We apply guided decoding Willard and Louf (2023) to ensure a machine-readable JSON format. These structured outputs are mapped to corresponding ICD codes using text-based matching. We investigate three distinct approaches: Zero-Shot, Few-Shot, and Zero-Shot with CoT prompting.
@@
In the Zero-Shot setting, we input the admission notes with task-specific instructions. Thus, the model can only leverage knowledge acquired from its initial pretraining and the information available in the shown example.
@@
Generative LLMs can learn from few-shot examples provided in the prompt Brown et al. (2020); Song et al. (2023); Parnami and Lee (2022) without adapting the parameters (in-context-learning). For this adaptation, either the user or a system incorporates complementary knowledge in the prompt. This is especially useful for domain specific tasks that are not well represented in the pretraining data.
@@ 4.1 Encoder
We fine-tune the pretrained encoders on the train split and select the best performing using AUROC on the validation split using AdamW Loshchilov and Hutter (2019) as optimizer and a polynomial learning rate decay. For the hyperparameter optimization, we apply Bayesian optimization. For more details compare Table˜4 and Table˜7.
Threshold tuning.
We optionally calibrate encoder outputs via class-wise threshold tuning on the validation split. For each label $c$, we evaluate $F_{1}$ over the sorted candidate threshold list $T_{c}$ and choose $j=\arg\max F_{c}$; we set $\tau_{c}=(T_{c}[j-1]+T_{c}[j])/2$ (or $\tau_{c}=T_{c}[0]$ if $j=0$). This midpoint rule avoids a fixed offset when thresholds are unevenly spaced. After calibration we still select the top-$n$ labels by score; tuning only changes label ordering, primarily affecting low-frequency labels.
@@ 4.2 Generative LLMs
We evaluate open-weight generative LLMs that are capable of processing sequences exceeding 4096 tokens. This extended sequence capacity is crucial for evaluating models in the few-shot setting, as sequence length increases with each added demonstration. Consequently, we excluded both available medical LLMs MedAlpaca and Med42 from our analysis, as both are based on the Llama2 architecture, which limits their maximum sequence length to 2048 tokens. We selected the following models for evaluation: Llama 3.1 8B and 70B, Mistral 7B, Mistral Nemo (13B), Qwen2 7B and 72B. For each model, we evaluate both the instruction fine-tuned and the non-instruction-tuned variants, resulting in a total of 12 models. We explicitly do not benchmark closed-source commercial models due to constraints of the Physionet license  [^2] https://physionet.org/content/mimiciv/view-license/2.2/.
@@
While generative models often produce valid JSON outputs and the correct number of prompted diagnoses, some of the evaluated models, especially those not fine-tuned with instruction data, frequently fail to do so. This limitation reduces overall task performance compared to encoder-based models. We report the performance of generating valid JSON output for all models in the Appendix in Table˜8. To mitigate this issue, we apply guided decoding Willard and Louf (2023) to enforce JSON output that contains exactly 20 strings. Each string represents a diagnosis description between 3 and 70 characters in length. To optimize performance, we limit the number of generated tokens to a maximum of 1500. Finally, we use greedy sampling with a temperature of 0.
@@
We use multiple prompts for each experiment: Zero-shot, zero-shot with CoT, and few-shot learning. The prompts consist each of a role description for the model and the respective task, further detailed in Section˜A.1.
@@
Because LLMs often struggle to predict exact ICD codes Boyle et al. (2023), we instead generate diagnostic descriptions and map them to ICD codes via their corresponding long or short descriptions. Based on a comparison of different approaches, we adopt NeuML-PubMedBERT [^3] https://huggingface.co/NeuML/pubmedbert-base-embeddings for text-to-label mapping. Details are reported in the Appendix in Table˜9. Finally, we truncate ICD codes to three digits to reduce label space and remove duplicates from the predictions.
@@ 4.3 Retrieval Models and Few-Shot Learning
We implement few-shot learning by incorporating one, three, or five demonstrations into the prompt. The effectiveness of in-context-learning depends on the selection of demonstrations Liu et al. (2022). Thus, we use a retriever to find the most relevant candidates. The retriever takes one admission note as input query and returns the most similar admission notes from the training split. We conduct three different retrieval strategies and compare them to random and gold demonstrations. Namely, we use document-based semantic similarity, latent outcome similarity (classifier fine-tuned on diagnosis prediction task), and a term-based retrieving strategy.
@@
To assess retrieval model performance for clinical outcome prediction, we evaluate them directly on the diagnosis prediction task. We apply majority voting over five retrieved training documents, selecting the top 20 most frequent labels, prioritizing documents with higher similarity in case of ties. We evaluate several semantic similarity retrievers based on S-BERT Reimers and Gurevych (2019) (see Section˜A.4 for complete list), selecting only the best performer for few-shot experiments to reduce computational costs. For latent outcome similarity retrieval, we use our fine-tuned BioMedBERT encoder (Section˜4.1), hypothesizing that admission notes with similar outcomes cluster in embedding space. We also employ BM25 Robertson and Zaragoza (2009) for term-based retrieval.
@@
4.4 Evaluation Protocol and Metrics
Fixed-size top-$n$ evaluation.
We evaluate all methods in a fixed-size top-$n$ setting, where each model outputs exactly $n$ diagnosis codes per admission note. A fixed output size is required to ensure comparability between encoder-based classifiers, which assign a score to each label, and generative LLMs, which produce an explicit list of codes.
For encoder models, we rank labels by predicted score and select the top-$n$ codes; when threshold tuning is applied, it affects score calibration and thus the ranking prior to top-$n$ selection. For generative models, we constrain decoding to return a JSON list containing exactly $n$ codes, thus simplifying guided decoding schema and reducing the maximum number of generated tokens per instance.
We choose n=20n{=}20 as a single global output budget that yields stable comparisons across partitions and supports a recall-oriented operating point. This value is motivated by the observed label cardinality in CliniBench: across partitions, notes contain on average 10.0–18.4 short codes (Table 1), and our preliminary analysis indicates an $F_{1}$ plateau around this range when varying $n$. We therefore use n=20n{=}20 to avoid systematic under-prediction on high-cardinality notes while keeping the output budget fixed across model classes. We treat $n$ as a benchmark hyperparameter that can be adjusted for different precision–recall preferences.
Evaluation Metrics.
We report macro-averaged recall, precision, and $F_{1}$ as main metrics. In addition, we report mean average precision (MAP), computed over the ranked list of predicted codes. Finally, we report main diagnosis accuracy (MD Acc.), i.e., whether the first annotated diagnosis code is contained in the predicted top-$n$ list.
Upper ceiling metrics.
A fixed top-$n$ protocol induces metric ceilings: for notes with fewer than $n$ codes, precision cannot reach 1.0; for notes with more than $n$ codes, recall cannot reach 1.0. These ceilings depend only on the set size $|Y|$, with $\mathrm{Prec}_{\max}(Y)=\min(|Y|,n)/n$ and $\mathrm{Rec}_{\max}(Y)=\min(|Y|,n)/|Y|$. On the test split with n=20n{=}20, the resulting dataset-level ceilings are Prec. =0.5566=0.5566 and Rec. =0.9824=0.9824.